OpenStudy (anonymous):

Identify the property shown. (7 • p) • 2k = 7 • (p • 2k)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

A. distributive B. associative C. commutative D. transitive

OpenStudy (skullpatrol):

When you multiply any three numbers does the grouping affect the result?

OpenStudy (anonymous):

I don't understand

OpenStudy (skullpatrol):

Do you know what grouping means?

OpenStudy (anonymous):

a set of people acting together with a common interest or purpose, esp. within a larger organization.

OpenStudy (skullpatrol):

Yes, this idea can be used with numbers too like (2*3)4=2(3*4)

OpenStudy (anonymous):

so it would be associative

OpenStudy (skullpatrol):

Yes, the grouping or association of the numbers does not affect the result.

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):

Thanks for trying to learn something new in math.

OpenStudy (anonymous):

@live10000000 @UsArmy3947 @

OpenStudy (anonymous):

plz stop tagging me I'm doing school work

OpenStudy (anonymous):

@ashley_f97 @Ashley1998 @Bella♥ @baby456

OpenStudy (ashley_f97):

Not trying to be rude or anything up you're question was answered. Why are you still tagging people?

OpenStudy (ashley_f97):

but*

OpenStudy (anonymous):

can u help me on this question please. Which statement shows the commutative property of multiplication? A. 7 · (4 · x) = (7 · 4) · x B. 18 + 6y = 6y + 18 C. If 5h = 30, then 30 = 5h D. (d + 4) ·17 = 17 · (d + 4)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

What is the commutative property of multiplication?

OpenStudy (anonymous):

A property of real numbers that states that the product of two factors is unaffected by the order in which they are multiplied; i.e., the product remains the same. Examples: 3 * 5 = 5 * 3 or 5 * x = x * 5.

OpenStudy (skullpatrol):

$$\huge ab=ba$$

OpenStudy (anonymous):

what is that

OpenStudy (skullpatrol):

That is the commutative property of multiplication.

OpenStudy (anonymous):

oh ok

OpenStudy (skullpatrol):